HUBBLE SPACE TELESCOPE DAILY REPORT #2051 PERIOD COVERED: 0000Z (UTC) 01/30/98 - 0000Z (UTC) 02/02/98 Daily Status Report as of 033/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/MA1/MA2 7920 (STIS MAMA Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make parallel field, ultraviolet imaging to search for Lyman-break galaxies at z>2.0, and z>1.0. The observations were completed as planned, and no problems were reported. 1.2 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at its nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and gain = 4 on each. The observations were executed as scheduled with no noted anomalies. 1.3 Completed NIC/3 7815 (NICMOS Earth Flats) The Near Infrared Camera (Camera Three) was used to make a Earth flat field calibration image. This proposal obtains a series of earth flats for a subset of narrow and medium band NIC/3 filters in advance of the NIC/3 campaign. These flats will be used to generate the master flats for narrow-band filters and to determine the OTA illumination pattern as a function of wavelength. The observations were completed as planned, and no anomalies were reported. 1.4 Completed Seven Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed three Sets of NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies NGC-6684, NGC-4192 and NGC-628 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.6 Completed NIC/1/2/3 7217 (GRISM Survey of Three Young Clusters) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the young cluster called the Trapezium. Although photometric surveys have been used to great success in identifying members of embedded star clusters, significant uncertainty surrounds the mapping of broadband photometry with masses and ages of the stars. The observations were completed as planned, and no problems were reported. 1.7 Completed NIC/1/2/3 7322 (The IMF Below 1 Msun in Young Clusters: 1.9 Micron Water Band Imaging of IC348) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of cluster J1-4423. This study will be the first to obtain spectral types for a young cluster population complete to K=16, and thereby quantify, to unprecedentedly high precision, the IMF below 1M{sun} in young clusters. The observations were executed as scheduled, with no anomalies. 1.8 Completed Three Sets of NIC/1/2/3 7235 (Observations of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the Hubble Deep Field area HDF-123645. This program observes a portion of the Hubble Deep Field with all three NICMOS cameras in the F160W and F110W filters to produce infrared images equivalent in depth to the existing WF/PC-2 HDF images. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.10 Completed NIC/1 7693 (NICMOS Pupil Transfer Function) The Near Infrared Camera (Camera One) was used to make observations of P177D in order to perform a pupil transfer function observations. The output of this program will be a two-dimensional surface fit that will give the large scale corrections needed to flat-field the images to an accuracy of about 2%. Also, changes in the PSF will be measured as a function of position within the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/1/2/3 7474 (Searching for Low Mass Stars: the Mass Function at the Hydrogen Burning Limit) The Near Infrared Camera (Cameras One, Two, and Three) was used to use the unprecedented IR imaging capabilities of NICMOS to study the low-mass stars at the end of the main sequence in the nearby globular cluster NGC-6121 spanning a wide range of metallicity, precisely at the wavelength corresponding to their maximum emission. These observations would complete our ongoing investigation by making it possible, for the first time, to reliably count stars right down to and beyond the main sequence H-burning limit, thus leaving no doubt as to whether the IMF indeed drops there or starts rising once again close to the limit. The observation completed as planned with no problems. 1.13 Completed Two Sets of WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.14 Completed NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the protoplanetary disk around star DK-Tau to search for water in its disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. As reported in 2.1, the initial acquisition for this proposal defaulted to fine lock on one FGS, and one observation may have been affected. Otherwise, the observations were completed as scheduled, with no anomalies. 1.15 Completed NIC/1/2/3 7608 (NICMOS Focus Monitoring) The Near Infrared Cameras (Cameras One, Two, and Three) were used to determine and monitor the optimal focus and tilt settings for all three NICMOS cameras by making observations of cluster NGC-3603. This proposal is built to run NIC1/2/3 sweeps on a 2-week basis. The observations completed without incident with no reported anomalies. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 23 Successful : 23 The acquisition scheduled at 032/061207Z defaulted to fine lock on FGS-1 only after FGS-3 had a scan step limit exceeded flag twice. HSTAR 6332 was written. The NIC/1/2/3 proposal described in 1.14 may have been affected. Scheduled Reacquisitions : 24 Successful : 24 2.2 FHST Updates: Scheduled : 42 Successful : 42 2.3 Operations Notes: This weekend, the NICMOS Hubble Deep Field (HDF) proposals executed nominally, although the NIC/1/2/3 proposal 7456 (described in 1.14) did have an initial acquisition that defaulted to fine lock on one FGS. One observation may have been affected. The completion of Sunday's activities marks the end of the three-week HDF campaign. The ground master image was updated at 032/0701Z. Secondary mirror commanding was successfully performed at 032/1644Z. Per HSTAR 6334, at 033/010556Z a NICMOS EDAC error was reported. No other error flags or status buffer messages were received. This NICMOS internal log error was reset at 033/0433Z per ROP NS-11. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH